the simple pendulum has a period of 2 seconds taking g=9.8ms-2.Calculate the length of the pendulum

Grade:11

1 Answers

Arun
4 years ago
 

Use T = 2π √( l / g )

Plug in the values, (you can square both sides for easier calculations & eliminating the square root though)

4 = 4 \Pi^{2} \frac{L}{g}

L = 9.8 / 9.8596 = 0.9939

Hope this helps